OpenStudy (darlene):

Write and equation for the line. use slope-intercept form, if possible. Through (5, 5), m=-2/9. Now i know the slope intercept formula is y=mx+b, and i also know that my slope is -2/9. now i am lost, i think

OpenStudy (anonymous):

y=-2/9x+b

OpenStudy (anonymous):

5=-2/9(5) +b

OpenStudy (dumbcow):

use the point (5,5) as your x,y values in the slope-intercept equation

OpenStudy (anonymous):

b=45/9+10/9 =55/9

OpenStudy (anonymous):

y=-2/9 x +55/9

OpenStudy (dumbcow):

also you can use point-slope form \[y-y1 = m(x-x1)\] \[y-5 = -(2/9)(x-5)\]